OpenStudy (anonymous):

ANSWER IN FIVE MINUTES FOR FAN + MEDAL!!! Oliver earns $7.50 per hour mowing lawns. He spends $9 for gas each week and $2 per week for lawnmower maintenance. If he mowed lawns for two and a half weeks in August for 15 hours each week, how much money did he earn in August? Use the expression w(7.5h - 9 - 2), where w represents the number of weeks he worked and h represents the number of hours he worked.

OpenStudy (k_lynn):

well first of all, how many weeks are there in August?

OpenStudy (anonymous):

There is a better expression that I can think of, but I will use the one it provides. Give me a minute.

OpenStudy (anonymous):

4 weeks per month.

OpenStudy (anonymous):

the answer is 281.25 dollars

OpenStudy (anonymous):

What does the problem look like when it's broken down?

OpenStudy (anonymous):

Don't give the answer, you must explain it.

OpenStudy (anonymous):

Okay so they worked 2.5 weeks, so 2.5(7.5h - 9 - 2)

OpenStudy (anonymous):

i did 7.50*15=112.50 then 112.5*2=281.25

OpenStudy (anonymous):

2.5(7.5(15) - 9 - 2)

OpenStudy (anonymous):

then u minus the 9 and 2 so if he uses 9dollars for gas and 2 for repairs then it would be 259.25

OpenStudy (anonymous):

$253.75 is correct.

OpenStudy (anonymous):

i got 270.25

OpenStudy (anonymous):

Original equation is w(7.5h - 9 - 2) plus in weeks for 2.5(7.5h - 9 - 2) then plug in hours per week for 2.5(7.5(15) - 9 - 2) = 253.75

OpenStudy (anonymous):

Hope that helps^

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

I'm gonna close the question now.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Thanks everyone that answered!